OpenStudy (anonymous):

Mary filled a con. with 6 1/4 gallons f oragne jucie .If a full glass of jucie equals 1/8, gallons How many glasses of orange jucie does the container hold/ ? To how ever answer this question please expain it greatly plz thx :)

OpenStudy (anonymous):

Just divide total vol given by vol of one glass

OpenStudy (anonymous):

? umm wat vol

OpenStudy (anonymous):

No of glasses of orange juice that the container holds = (6 1/4) / (1/8) 25 8 200 = (25/4) / (1/8) = ---- x ---- = ----- = 50 glasses 4 1 4

OpenStudy (anonymous):

Hope this is clear.....

OpenStudy (anonymous):

yes act. its is thx :)

OpenStudy (anonymous):

U r welcome anytime...:)

OpenStudy (anonymous):

ok i have aquestion were did yuh gt the 25/4 from?